can you please help me solve the problems thanks 1. 90 miles over 1 hr equals feet over seconds
2. 3.5 cm over 1 second equals meters over 1 min
3. 114 over 1 yr equals liters over 1 day
4. 500 dallors over 8 hrs equal cents over 1 day

Answer by stanbon(75887)

1. 90 miles over 1 hr equals feet over seconds
(90 mi/1 hr)
= (90*5280 ft/3600 sec)
= (475200/3600) ft/sec
= 132 ft/sec
----------------
2. 3.5 cm over 1 second equals meters over 1 min
(3.5 cm/1 second)(60 seconds/1 min)(1 meter/100 cm)
= (3.5*60*10/(1*1*100)(meters/minute)
= 21 (meters/ minute)
-----------------
3. 114??? over 1 yr equals liters over 1 day
-------------------
4. 500 dollars over 8 hrs equal cents over 1 day
(500 dollars/8 hrs)
= (500*100 cents/(8/24) day)
= (50000/(1/3)) cents/day
= 150000 (cents/day)
===============
Cheers,
Stan H.